OpenStudy (anonymous):

determine the equation of the line through the given points: (-3,5), (1,7)

OpenStudy (anonymous):

If a particular point on a line and the slope of a line are known, we can write an equation for that line, using Point-Slope Form: An equation of the line with slope m passing through the point is y-y1=m (x-x1) must first find the slope of the line, using the formula for slope: m=(y2-y1)/(x2-x1) or m=(7-5)/1-(-3)) .... finish it

OpenStudy (anonymous):

kay, got the answer thank you so much!